OpenStudy (anonymous):

Find the standard form of the equation of the hyperbola with the given characteristics. foci: (+-3,0) Asymptotes: y=+-2x

OpenStudy (amistre64):

since the foci are located parallel to the x axis; we can make a note of that and say that they open to the left and right

OpenStudy (anonymous):

I know the origin is at (0,0) my equation should look like: y^2/a^2-x^2/b^2=1

OpenStudy (amistre64):

the center appears to already be at 0; otherwise youd have different 'x' parts for the foci

OpenStudy (anonymous):

jk it would be x^2/a^2-y^2/b^2=1

OpenStudy (amistre64):

the asymptote tells us what a and b are really; since it is in the form y = (b/a)x

OpenStudy (amistre64):

2 b -- x = --- x 1 a the y component is b and the x compnent is a

OpenStudy (anonymous):

Right and I know I need to find c^2=a^2+b^2 I know C is 3^2 right?

OpenStudy (amistre64):

x^2 y^2 --- - --- = 1 1^2 2^2

OpenStudy (amistre64):

a is only needed if you are looking to find a or b; we know both a and b from the asympotote

OpenStudy (anonymous):

ok but that can't be right....I think we may have a wrong....

OpenStudy (anonymous):

my options are x^2 y^2 --- - --- = 1 9 4 x^2 y^2 --- - --- = 1 9/5 36/5 x^2 y^2 --- - --- = 1 36/5 9/5

OpenStudy (anonymous):

x^2 y^2 --- - --- = 1 9 4 I would guess this one but I'm not sure

OpenStudy (amistre64):

lets recheck then :)

OpenStudy (amistre64):

we need to determine the "rectangle" of the hyperbola; to so that, lets write it as an ellipst

OpenStudy (anonymous):

uh.. ok?

OpenStudy (amistre64):

x^2 y^2 --- + ---- = 1 the vertices of this ellipse are the top and sides a^2 b^2 of your asympotoe information

OpenStudy (anonymous):

correct

OpenStudy (amistre64):

2x is the asympy.... this tells us the from the center, we go up 2 and over 1 to for the top right corner of the rectangle; and our ellipse will be 2 high and 1 left right?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

ok I got it

OpenStudy (amistre64):

and it should end up looking like this inthe end

OpenStudy (anonymous):

ok.... ugh this stupid math class their answer key never matches up to what I get lol

OpenStudy (amistre64):

our y compnent is 2^2 regardless right :)

OpenStudy (amistre64):

id go with your first option; dbl chk by making y=0 and seeing what you get for an x value.. maybe :)

OpenStudy (anonymous):

ok thanks